OpenStudy (anonymous):

(7u+6v)(7u-6v) how can u do this?

OpenStudy (saifoo.khan):

Difference of squares.

OpenStudy (akshay_budhkar):

\[(a+b)(a-b)=a^2-b^2\]

OpenStudy (anonymous):

You can FOIL it

OpenStudy (akshay_budhkar):

\[a=7u , b=6v\]

OpenStudy (anonymous):

7u*7u=49u^2 7u*6v and 7u*-6v cancel out 6v*-6v=-36v^2 49u^2-36v^2

OpenStudy (anonymous):

so 49u^2-36v^2 will be the answer cuz i get something different? =(

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

thanks very much =)